Given the following information, compute the missing columns in the life table:

Age (x) Number Alivebx Ix SX mx 452 334 122 75 0

After completing the table, compute the number of individuals in each age class for the next two generations.
Answer #1
X ax bx lx sx mx
0 452 0 1 0.74 0
1 334 2 0.74 0.37 2
2 122 3 0.27 0.61 3
3 75 1 0.17 0.29 1
4 22 1 0.05 0 1
5 0 0 0 - 0

Sample calculations

lx = Nx/No = 122/452 = 0.27

Sx = Nx+1/Nx = 122/334 = 0.37

Mx = bx According to Ricklefs text.